Two sides of a parallelogram are 120 degrees with each other. The lengths of the sides

are 10.0 m and 6.5 m. Find the length of the diagonal opposite the included angle.

3. The base of the triangle is 5 m and the base angles are 65 and 40. Find the other side of

the triangle.


1
Expert's answer
2021-08-30T15:05:19-0400

Solution.

"\alpha=120^o;"

"a=10.0m;"

"b=6.5m;"

"c^2=a^2+b^2-2abcos\alpha" ;

"c=\sqrt{100+42.25+2\sdot10\sdot6.5\sdot0.5}=14.4m;"

"3. a=5m;"

"\beta=65^o;"

"\gamma=40^o;"

"\dfrac{a}{sin\alpha}=\dfrac{b}{sin\beta}\implies b=\dfrac{asin\beta}{sin\alpha};"

"b=\dfrac{5\sdot0.9}{0.97}=4.6m;"

"\dfrac{a}{sin\alpha}=\dfrac{c}{sin\gamma}\implies c=\dfrac{asin\gamma}{sin\alpha};"

"c=\dfrac{5\sdot0.64}{0.97}=3.3m;"

Answer:"c=14.4m;"

"3. b=4.6m; c=3.3m."